### v1

# Stabilization of active tissue deformation by a dynamic morphogen gradient

Abstract
A key process during animal morphogenesis is oriented tissue deformation, which is often driven by internally generated active stresses. Yet, such active oriented materials are prone to well-known instabilities, raising the question of how oriented tissue deformation can be robust during morphogenesis. In a simple scenario, we recently showed that active oriented deformation can be stabilized by the boundary-imposed gradient of a scalar field, which represents, e.g., a morphogen gradient in a developing embryo. Here, we discuss a more realistic scenario, where the morphogen is produced by a localized source region, diffuses across the tissue, and degrades. Consistent with our earlier results, we find that oriented tissue deformation is stable in the gradient-extensile case, i.e. when active stresses act to extend the tissue along the direction of the gradient, but it is unstable in the gradient-contractile case. In addition, we now show that gradient-contractile tissues can not be stabilized even by morphogen diffusion. Finally, we point out the existence of an additional instability, which results from the interplay of tissue shear and morphogen diffusion. Our theoretical results explain the lack of gradient-contractile tissues in the biological literature, suggesting that the active matter instability acts as an evolutionary selection criterion.
I Introduction
Oriented tissue deformation is a key process during animal morphogenesis [1], including for instance body axis elongation [2, 3, 4] and organ formation [5, 6, 7, 8, 9, 10]. In many cases, such tissue deformation is at least in part driven by internally generated active stresses [11, 12, 13, 8, 10], and in recent years, deforming tissues have successfully been described as active materials [14, 8, 15, 16, 17, 18, 19, 20, 21]. Moreover, to achieve persistent large-scale oriented deformation, tissues need some kind of tissue-wide orientational information, which can be represented for instance by a polar or a nematic field [8, 15, 22, 23, 19]. However, active polar or nematic materials are subject to the well-known Simha-Ramaswamy instability, where an initially globally ordered orientational field loses this order due to active flows [24, 25, 26]. This raises the question of how oriented tissue deformation can be stable during animal development.
In animals, oriented tissue deformation can be affected by tissue-scale concentration patterns of proteins, which are often called morphogens. Such morphogen patterns are known to coordinate developmental processes on the tissue or even organism scale [1]. In many cases, they provide positional information [27, 3, 28]. Yet, in some cases they also provide orientational information that defines the axis of active stresses driving oriented tissue deformation [5, 29, 6, 12, 7], typically by guiding cellular polarity and cell rearrangements [2, 6, 12, 4, 30]. Such alignment of orientational information with morphogen gradients has recently also been discussed theoretically [31].
In a recent publication [32], we theoretically showed that oriented tissue deformation can be stabilized by morphogen gradients. We used a hydrodynamic model for an active polar material and showed that the gradient of a scalar field, which describes the morphogen gradient, can stabilize the otherwise unstable material. If the scalar field advects with the material flows, we found that the stability of oriented material deformation depends on the coupling between scalar field gradient direction and deformation axis. In the gradient-extensile case, i.e. when anisotropic active stresses act to extend the material along the gradient direction, the scalar field can stabilize the active material deformation. However, in the gradient-contractile case, i.e. when anisotropic active stresses act to contract the material along the gradient direction, the deformation is unstable. However, one could always stabilize even a gradient-contractile system by sufficiently increasing morphogen diffusion.
In our past publication [32], we have created a uniform scalar field gradient in a simple way through appropriate boundary conditions. While providing advantages for analytical calculations, this is quite different from the real mechanism creating morphogen gradients in developing biological tissues [33, 34, 1, 35, 36]. Morphogens do indeed advect with tissue flows [37], which we already considered in our previous work [32]. However, they are typically produced by some source region, diffuse across the tissue, and degrade [38, 36]. As a consequence, depending on the source region, this can lead to approximately exponentially decaying [33, 39, 40, 35] or periodic [41, 42] concentration profiles. However, it is not known how such morphogen dynamics would affect oriented deformation of an active material.
In this article, we study the stability of active material deformation when accounting for more realistic morphogen dynamics (Figure 1A). We consider an active viscous material with local flow velocity $\bm{{\mathrm{v}}}(\bm{r},t)$, where active stresses are defined by the gradient of a morphogen field $c(\bm{r},t)$. We extend our previous work by more realistic morphogen dynamics. Specifically, the morphogen field $c(\bm{r},t)$ is secreted by a source region that is defined by a field $s(\bm{r},t)$, which is conserved and only advects with tissue flows. The morphogen field $c(\bm{r},t)$ not only advects with tissue flows, but also diffuses and degrades at constant rates everywhere.
Using linear stability analysis, we find that also when including more realistic morphogen dynamics, gradient-extensile materials are stable, while gradient-contractile materials are unstable. Yet additionally, we also show that in contrast to a boundary-provided morphogen gradient [32], in the case of a secreted morphogen gradient, a gradient-contractile system can never be stabilized by diffusion. Finally, we find an additional instability in deforming systems for both gradient-extensile and gradient-contractile cases, which arises from the interaction between shear and diffusion, and which is unrelated to the Simha-Ramaswamy active matter instability.
In the following, we first introduce the model in section II. We then study the case of a uniform gradient in both $s(\bm{r},t)$ and $c(\bm{r},t)$ for fixed system size in section III, for which the linear stability can be computed analytically. Next, in section IV, still for fixed system size, we study the more realistic case of a localized source $s(\bm{r},t)$, which gives rise to a non-uniform morphogen gradient $c(\bm{r},t)$. We then discuss the stability of freely deforming systems in section V. Finally, in section VI, we discuss our findings in the context of the biological literature.
Figure 1: (A) Hydrodynamic model. A spatially varying source field $s(\bm{r},t)$ secretes a morphogen with concentration $c(\bm{r},t)$, which also diffuses and degrades. The gradient of $c$ creates extensile/contractile active anisotropic stresses, which drive viscous flows with velocity $\bm{{\mathrm{v}}}(\bm{r},t)$. These flows affect both source and morphogen fields through advection. (B) Typical orders of magnitude for tissue deformation rates (top), morphogen diffusion rates (middle), and morphogen degradation rates (bottom) from the biological literature. Blue marks are data from the fruit fly Drosophila melanogaster, and orange marks are data from zebrafish, Danio rerio. Tissue deformation rates in the fruit fly were estimated from: egg chamber [43, 44], dorsal thorax [12], pupal wing [8], hindgut [45, 5], renal tubules [7], and germ-band [11]; and in zebrafish from mesodermal explants [46] and the tailbud [47]. Morphogens: Decapentaplegic in the fruit fly [48, 33, 36]; Cyclops, Squint, Lefty1 and Lefty2 in zebrafish [34, 39].
II Model
II.1 Hydrodynamic fields and bulk dynamics
We describe the tissue as a 2D active material, where we consider three interacting hydrodynamic fields: a scalar field $s(\bm{r},t)$, which describes the source, i.e. cells in the tissue which secrete the morphogen; a scalar field $c(\bm{r},t)$, which describes the morphogen concentration; and a vector field $\bm{{\mathrm{v}}}(\bm{r},t)$, which describes the tissue flows (Figure 1A). We find that literature values for the respective rates of tissue deformation, morphogen diffusion, and morphogen degradation are of roughly similar orders of magnitude (Figure 1B), suggesting that none of them can be easily neglected, and we thus include all three effects in our model.
The source field $s(\bm{r},t)$, which describes morphogen production, only advects with material flows:
$\frac{\partial s}{\partial t}+\partial_{i}({\mathrm{v}}_{i}s)=0,$ (1)
where $\partial_{i}$ denotes the partial derivative with respect to the spatial coordinate $r_{i}\in\{x,y\}$. Here and in the following, we label spatial dimensions by Latin indices $i,j,\dots$ and adopt the Einstein notation of summing over repeated indices.
The morphogen field $c(\bm{r},t)$ follows a generalized conservation equation:
$\frac{\partial c}{\partial t}+\partial_{i}({\mathrm{v}}_{i}c)=k_{\mathrm{p}}s-k_{\mathrm{d}}c+D\partial_{i}^{2}c.$ (2)
Here, we have introduced production rate $k_{\mathrm{p}}$, degradation rate $k_{\mathrm{d}}$, and diffusion coefficient $D$.
Tissue flows $\bm{{\mathrm{v}}}(\bm{r},t)$ are governed by overdamped, incompressible, viscous dynamics:
$\partial_{i}\sigma_{ij}$ $=0,$ (3)
$\partial_{i}{\mathrm{v}}_{i}$ $=0,$ (4)
where the stress tensor is given by
$\sigma_{ij}=2\eta\tilde{v}_{ij}-\Pi\delta_{ij}+\tilde{\sigma}^{a}_{ij}.$ (5)
Here, $\eta$ is the shear viscosity, $\tilde{v}_{ij}:=(\partial_{i}{\mathrm{v}}_{j}+\partial_{j}{\mathrm{v}}_{i}-\partial_{k}{\mathrm{v}}_{k}\delta_{ij})/2$ is the shear rate tensor, $\Pi$ is the hydrostatic pressure, and we use the following expression for the active stress $\tilde{\sigma}^{a}_{ij}$:
$\tilde{\sigma}^{a}_{ij}=\alpha\left[(\partial_{i}c)(\partial_{j}c)-\frac{1}{2}(\partial_{k}c)^{2}\delta_{ij}\right].$ (6)
This expression for the anisotropic active stress is the same as in Active Model H [49, 50]. Its direction is determined by the sign of the coefficient $\alpha$: For $\alpha<0$, the activity is gradient-extensile, i.e. it acts to extend the material along the local $c$ gradient direction and contract perpendicularly; while for $\alpha>0$, the activity is gradient-contractile, i.e. it acts to contract the material along the $c$ gradient direction and extend perpendicularly (Figure 1A).
The expression for the active stress, Eq. (6), is motivated by what is known about developing biological tissues. Specifically, anisotropic internal stresses are usually generated by an anisotropic distribution of cytoskeletal elements within cells. Such an anisotropic distribution is created by cell polarity, which in turn can be controlled by morphogen gradients [12, 51, 52, 7, 30]. Here, we assume that the time scales involved in these different couplings are fast enough, so that we directly couple the generated active stresses to the morphogen gradient. Eq. (6) represents the lowest-order term for such a coupling [49].
II.2 Boundary Conditions
We use periodic boundary conditions, where the periodic box has generally time-dependent dimensions $L_{x}(t)\times L_{y}(t)$ and constant area, $L_{x}(t)L_{y}(t)=\mathrm{const}$. Throughout this article, we always use one of the following two boundary conditions:
  1. 1.
    We fix the system dimensions $L_{x}$ and $L_{y}$, or
  2. 2.
    we allow the system to freely deform by fixing the externally applied stress to zero, which results in a pure shear deformation of the periodic box. Integration of Eq. (5) implies that the instantaneous box shear rate is
    $\frac{1}{L_{x}}\,\frac{\mathrm{d}L_{x}}{\mathrm{d}t}=-\frac{\alpha}{4\eta}\int{\big{[}(\partial_{x}c)^{2}-(\partial_{y}c)^{2}\big{]}\;\mathrm{d}A},$ (7)
    where the integral is over the whole periodic box.
We apply fixed system dimensions in section III and section IV, and study a freely deforming system in section V.
We investigate the stability of two stationary states: (i) a state where a linear source profile $s$ gives rise to a linear morphogen profile $c$, and (ii) a state where a localized source $s$ generates a spatially decaying morphogen field $c$. In both cases, the amplitude of the source field is defined by a constant $s_{b}$. In case (i), which we study in section III, the boundary conditions are periodic with the exception of the vertical boundaries for the scalar fields $s$ and $c$, where we set for all $y\in[0,L_{y})$:
$s(0,y)$ $=s(L_{x},y)-s_{b},$ (8)
$c(0,y)$ $=c(L_{x},y)-\frac{k_{\mathrm{p}}}{k_{\mathrm{d}}}s_{b},$ (9)
with fixed $s_{b}$. These modified conditions remove the discontinuity that would otherwise appear for linear profiles in $s$ and $c$ at the boundary, and thus ensure that linear $s$ and $c$ profiles can be stationary [32]. In case (ii), which we study in section IV and section V, we use standard periodic boundary conditions without offset.
II.3 Dimensionless equations
We nondimensionalize the set of equations Eqs. (1)–(9) in the following way:
  • We choose the unit of length such that the system area is one $L_{x}L_{y}=1$.
  • We choose the unit of the morphogen field $c$ such that $k_{\mathrm{p}}=k_{\mathrm{d}}$.
  • We choose the units of time and of the source field $s$, such that the free deformation rate of a squared system is 1. To this end, we set $|\alpha|/4\eta=1$, and in case (i) of a linear gradient, we also set $s_{b}=1$, while in case (ii) of a non-uniform gradient, we fix $s_{b}$ as described in section IV (Appendix E.1).
The last condition allows to compare the linear stability of systems with different source profiles, beacuse the reference time scale is always the respective free deformation rate.
Rescaled accordingly, the dimensionless equations read:
$\frac{\partial s}{\partial t}+{\mathrm{v}}_{i}\partial_{i}s$ $=0,$ (10)
$\frac{\partial c}{\partial t}+{\mathrm{v}}_{i}\partial_{i}c$ $=\left(D\partial_{i}^{2}-k_{\mathrm{d}}\right)c+k_{\mathrm{d}}s,$ (11)
00 0 $=\frac{1}{4}\partial_{i}^{2}{\mathrm{v}}_{j}-\partial_{j}\Pi^{\prime}+\mathrm{sgn}(\alpha)\,\partial_{i}\big{[}(\partial_{i}c)(\partial_{j}c)\big{]},$ (12)
$\partial_{i}{\mathrm{v}}_{i}$ $=0.$ (13)
Here, we have introduced the sign function, $\mathrm{sgn}(\alpha):=\alpha/|\alpha|$, and have set $\Pi^{\prime}:=\Pi+\mathrm{sgn}(\alpha)(\partial_{k}c)^{2}/2$. The offset introduced at the boundary conditions in Eqs. (8) and (9) is rescaled to $1$ for both scalar fields.
III Linear source profile
To build intuition, we study for fixed box dimensions $L_{x}=L_{y}=1$ the linear stability of a stationary state with linear spatial profiles in $s$ and $c$ (Figure 2A), given by:
$s_{0}$ $=x,$ $c_{0}$ $=x,$ $\bm{{\mathrm{v}}}_{0}$ $=\bm{0}.$ (14)
There are no flows in the stationary state, because the gradient of $c$ is homogeneous, and thus the active stress is homogeneous across the system of fixed dimensions.
In this case, linear stability can be computed analytically. When we add a small perturbation around this state,
$s$ $=x+\delta s,$ $c$ $=x+\delta c,$ $\bm{{\mathrm{v}}}$ $=\bm{\delta{\mathrm{v}}},$ (15)
and linearize the dynamics, only constant prefactors occur. As a consequence the solutions in $(\delta s,\delta c,\bm{\delta{\mathrm{v}}})$ are spatial Fourier modes with wave vectors $\bm{q}=q(\cos{\phi},\sin{\phi})$, whose amplitudes grow or shrink exponentially with time $t$ (Appendix D).
Figure 2: Stability of a fixed-size system with a linear source profile: gradient-contractile systems (orange) are unstable, whereas gradient-extensile systems (green) are marginally stable. (A) The linear source profile, $s_{0}(x)=x$ in dimensionless units. In the stationary state, $c_{0}(x)=s_{0}(x)$. (B) The perturbation growth rates $\omega$ depending on wave vector angles $\phi$. There are 2 branches for each of the 4 cases reported (see legend). (C) Effect of morphogen diffusion rate $Dq_{0}^{2}$ (with $q_{0}=2\pi$) on the maximal perturbation growth rates $\omega_{\mathrm{max}}$ of the linear systems. Inset: same data in log-log scaling for the gradient-contractile case.
We first discuss the limit without diffusion, $D=0$. When we subtract the linearized versions of the source dynamics, Eq. (10), from the linearized version of the morphogen dynamics, (11), we obtain:
$\frac{\partial}{\partial t}\big{(}\delta c-\delta s\big{)}=-k_{\mathrm{d}}\big{(}\delta c-\delta s\big{)}.$ (16)
Thus, any difference between morphogen field $c$ and source field $s$ decays at rate $k_{\mathrm{d}}$ (dashed horizontal lines in Figure 2B). In addition, the combination of morphogen dynamics and generated active flows, Eqs. (11)–(13), leads to the same results as for a single scalar field discussed in earlier publications [50, 32] (Appendix D.1). This means that the perturbation growth rate $\omega_{\mathrm{act}}$ due to the active flows depends on the angle of the Fourier mode $\phi$ as $\omega_{\mathrm{act}}:=4\,\mathrm{sgn}(\alpha)\sin^{2}{\phi}$, as indicated by the dashed curves in Figure 2B. Thus, without diffusion, the system is marginally stable in the gradient-extensile case, where the maximal perturbation growth rate is zero (green dashed curve), and unstable in the gradient-contractile case, where the maximum perturbation growth rate is strictly positive (orange dashed curve).
Even in the presence of diffusion, $D>0$, a gradient-contractile system remains always unstable, although the maximal perturbation growth rate decreases with diffusion (orange curve in Figure 2C & inset). This is in contrast to earlier findings for a system with a single scalar field $c$ without source, where we found that diffusion can stabilize the system in the gradient-contractile case. To understand where the difference comes from, we consider the limit of very large $D$. The linearized version of the morphogen dynamics, Eq. (11), in Fourier space reads:
$\frac{\partial\delta c}{\partial t}+\delta{\mathrm{v}}_{x}=-\left(Dq^{2}+k_{\mathrm{d}}\right)\delta c+k_{\mathrm{d}}\delta s.$ (17)
The flow field $\bm{\delta{\mathrm{v}}}$ enters on the left-hand side due to advection, where $\partial_{x}c_{0}=1$ in dimensionless units. The flow $\bm{\delta{\mathrm{v}}}$ is actively created by the morphogen field itself through Eqs. (12) and (13), and their $x$ component in Fourier space reads: $\delta{\mathrm{v}}_{x}=-\omega_{\mathrm{act}}\delta c$, with the active flow time scale $\omega_{\mathrm{act}}$ introduced before (Appendix D). Denoting the minimal non-zero wave vector magnitude in our system by $q_{\mathrm{min}}$, we take the limit where $Dq_{\mathrm{min}}^{2}$ is much larger than all other relevant time scales in the system. In this case, according to Eq. (17), the Fourier amplitudes $\delta c$ for each wave vector relax adiabatically fast towards
$\delta c\simeq\frac{k_{\mathrm{d}}}{Dq^{2}}\delta s.$ (18)
This effectively corresponds to a system where we again just have a single scalar field $s$, which advects with active tissue flows $\bm{{\mathrm{v}}}$, but does not diffuse. The active tissue flows are generated by the perturbations of $c$, and their $x$ component is given by $\delta{\mathrm{v}}_{x}=-\omega_{\mathrm{act}}\delta c\simeq-(\omega_{\mathrm{act}}k_{\mathrm{d}}/Dq^{2})\delta s$. Inserting this in the linearized version of the source dynamics in Fourier space, we get:
$\frac{\partial\delta s}{\partial t}=-\delta{\mathrm{v}}_{x}=\frac{\omega_{\mathrm{act}}k_{\mathrm{d}}}{Dq^{2}}\delta s=\frac{4k_{\mathrm{d}}\mathrm{sgn}(\alpha)\sin^{2}{\phi}}{Dq^{2}}\delta s.$ (19)
Thus, the sign of the growth rate of $\delta s$ just depends on the sign of $\alpha$. As a consequence, the system is unstable for a gradient-contractile system ($\alpha>0$), even in the limit of a large diffusion constant. The reason for this is that the source field itself advects with the active flows, but does not diffuse. For the same reason, any gradient-extensile system remains only marginally stable even in the presence of morphogen diffusion.
IV Localized source profile
In the following, we discuss how the linear stability changes when the source profile is not linear but instead localized to some region of finite width $\mathrm{{w}}$. This is motivated by what we know of developing biological tissues [1, 38, 33, 42]. In this section, we discuss a system of fixed size, while in section V, we discuss a freely deforming system.
As initial state, we use a source profile that is similar to a von Mises distribution, which is a generalization of a Gaussian to periodic boundary conditions:
$s_{0}(\bm{r})=s_{b}\Bigg{[}\exp{\bigg{(}\frac{\cos{(q_{0}x)}}{(q_{0}\mathrm{{w}})^{2}}\bigg{)}}-\exp{\bigg{(}-\frac{1}{(q_{0}\mathrm{{w}})^{2}}\bigg{)}}\Bigg{]}$ (20)
with $q_{0}=2\pi/L_{x}$. Here, $\mathrm{{w}}$ is a parameter that adjusts the width of the source region, and the second term in Eq. (20) ensures that $s_{0}(\bm{r})$ is zero at its minimum. As initial $c_{0}$ field, we choose the $c$ field that is stationary in the absence of flows for $s_{0}$ given by Eq. (20). In Fourier space:
$c_{0}(\bm{q})=\frac{k_{\mathrm{d}}}{k_{\mathrm{d}}+Dq^{2}}s_{0}(\bm{q}).$ (21)
Thus, for $D=0$, morphogen and source profiles coincide, and examples for $s_{0}(x)=c_{0}(x)$ for different $\mathrm{{w}}$ values are shown in Figure 3A. Meanwhile, finite diffusion $D>0$ smoothens the morphogen profile $c_{0}(x)$ as compared to the source profile $s_{0}(x)$ (Figure 4A).
Figure 3: Stability of a fixed-size system with a localized source profile and a non-diffusive morphogen, $D=0$. (A) The localized source profile $s_{0}(x)$ from Eq. (20) for different values of the source width $\mathrm{{w}}$. (B,C) Growth rates $\omega_{m}$ of perturbation modes depending on their mean wavenumber $\langle n\rangle_{m}$ for a system with $\mathrm{w}=1/2\pi$, for (B) a gradient-contractile and a (C) gradient-extensile system. The eigenvalue problem was solved numerically as described in Appendix C.3. (D) Zoomed region of panel C displaying high-wavenumber perturbation modes with $\omega>0$, which we interpret as numerical artifacts, due to the spatial discretization (see main text, Appendix E.2, and panel G). (E-G) Morphohen field perturbation $\delta c(\bm{r})$ for three different eigen modes (as indicated in panels B and D). (E) $\delta c$ mode for small wavenumber $\langle n\rangle$, with positive growth rate in the gradient-contractile case. (F) Typical $\delta c$ mode with high $\langle n\rangle$ and high growth rate in the gradient-contractile case. (G) Typical $\delta c$ mode with numerically obtained positive growth rate in the gradient-extensile system. (H) Maximal growth rate $\omega_{\mathrm{max}}$ in gradient-contractile and -extensile systems depending on the numerical discretization, $N_{x}$. For gradient-extensile systems, this growth rate decreases as $\omega_{\mathrm{max}}\sim N_{x}^{-2}$. (I) Perturbation growth rate depending on the source width $\mathrm{{w}}$. Squares indicate $\omega_{\mathrm{max}}$, whereas triangles indicate the growth rate of perturbation modes in with small $\langle n\rangle$ (see main text and panel E).
In this initial state, there are no flows in the system, $\bm{{\mathrm{v}}}_{0}=\bm{0}$. This is because the active stress in this case is parallel to the $x$ axis and homogeneous along the $y$ axis, because of incompressibility, and of the fixed system size, which together imply that there are no Stokes flows (Appendix B.2).
For a localized source, we study linear stability numerically. This is because the non-uniform gradient in the $c$ field creates non-constant coefficients in the linearized perturbation dynamics. As a consequence, the linearized dynamics in Fourier space now includes Fourier-space convolutions of the stationary profile $c_{0}$ and the morphogen perturbations $\delta c$ (Appendix C.2). Because of this, while the eigen modes of the linearized dynamics still grow or shrink exponentially with time, they are not single Fourier modes any more. We thus compute the perturbation eigen modes and their respective growth rates numerically.
For the case without diffusion, $D=0$, we plot in Figure 3B, C the growth rates $\omega_{m}$ of the different eigen modes $m$, where Figure 3B is for a gradient-contractile system and Figure 3C for a gradient-extensile system. For each eigen mode, the growth rates are plotted over their respective mean wavenumber $\langle n\rangle_{m}$, which we define as
$\langle n\rangle_{m}:=\frac{1}{2\pi}\,\frac{\sum_{\bm{q}}{\big{|}\delta c_{m}(\bm{q})\big{|}\,q}}{\sum_{\bm{q}}{\big{|}\delta c_{m}(\bm{q})\big{|}}},$ (22)
in dimensionless units (i.e. $L_{x}=L_{y}=1$), and where $\delta c_{m}(\bm{q})$ denotes the morphogen field part of eigen mode $m$ in Fourier space. For the gradient-contractile case, Figure 3B, we thus find that also for a non-uniform source the system is unstable, since there are many eigen modes with a positive growth rate $\omega_{m}>0$.
Surprisingly, Figure 3C seems to suggest that also gradient-extensile systems are unstable in this case, because there are modes with a positive growth rate (see magnified region in Figure 3D). However, these modes all have almost maximal wave number $\langle n\rangle_{m}\gg 1$ (Figure 3D, G). We thus wondered whether these positive growth rates might a numerical artifact, for instance due to the numerical discretization of the source and morphogen fields. To test this, we quantified the maximum of all growth rates $\omega_{m}$ for increasing numerical resolution in Figure 3H. We find that for the gradient-extensile case (green squares) the maximal growth rate decreases as $\omega_{\mathrm{max}}\sim N_{x}^{-2}$ with the numerical resolution in $x$ direction, $N_{x}$, while it remains largely constant for the gradient-contractile case (orange squares). This indicates that the positive growth rates of the large-$\langle n\rangle_{m}$ modes in the gradient-extensile case shown in Figure 3C inset are due to a numerical discretization artifact, without which the maximal growth rate in the gradient-extensile case would be zero. We independently confirm this further in appendix Appendix E.2.
Next, we study for $D=0$ how the system’s stability depends on the source width $\mathrm{{w}}$ (Figure 3I). The gradient-contractile system (orange symbols) is always unstable, where the squares in Figure 3I show the maximum growth rate, which decreases with source size, but approaches a positive value for $\mathrm{{w}}\rightarrow\infty$. As indicated by Figure 3B, this maximal growth rate can correspond to a mode $m$ with a large wavenumber $\langle n\rangle_{m}$ (Figure 3F). However, in biological tissues there is a natural upper wave vector cutoff, provided for instance by the cell size. Thus, to test in how far our results are still valid in the presence of some wave vector cutoff, we also indicate the growth rates of a small-wavevector mode. We selected the mode with $\langle n_{x}\rangle_{m}\neq 0$ and $\langle n_{y}\rangle_{m}\neq 0$ that has the smallest $\langle n\rangle_{m}$, where we used $\langle n_{i}\rangle_{m}:=\left(\sum_{\bm{q}}{\big{|}\delta c_{m}(\bm{q})\big{|}\,q_{i}}\right)/\left(2\pi\sum_{\bm{q}}{\big{|}\delta c_{m}(\bm{q})\big{|}}\right)$. We find that also this small-wavevector mode has a positive growth rate for all source widths, which also approaches a finite value for $\mathrm{{w}}\rightarrow\infty$ (orange triangles in Figure 3E). Thus, we expect the material to be unstable even in the presence of some upper wave vecor cutoff. In contrast, the gradient-extensile system (green symbols) is always marginally stable, where the squares show again the maximum growth rate. The squares are open to indicate that we ignored any modes with large wavevecor and positive growth rates based on our discussion in the previous paragraph.
Figure 4: Effect of finite morphogen diffusion, $D>0$, on the stability of the fixed-size system with a localized source. Diffusion can not stabilize a gradient-contractile system. (A) Stationary morphogen profiles $c_{0}(x)$ for different diffusion coefficients, where $\mathrm{{w}}=1/2\pi$. (B) Perturbation growth rates of gradient-contractile systems for the three diffusion coefficients indicated in panel A. (C) Perturbation growth rates with increasing $D$. Squares indicate maximal growth rates $\omega_{\mathrm{max}}$ and triangles indicate growth rates of modes in the small $\langle n\rangle$ limit, $\omega_{\mathrm{s}}$.
Finally, we study how the system stability depends on the diffusion coefficient $D$. In this case, gradient-extensile systems are still marginally stable for any $D>0$ (Figure 4C). They are not further stabilized by diffusion for the same reason as for a linear source profile: any perturbation with $\phi=0$ in the source field neither induces any flows, nor can it “diffuse away”.
For gradient-contractile systems, diffusion decreases the growth rates of the large-wavevector modes (Figure 4B). Thus, for large enough diffusion constant, the system’s fastest growing mode becomes the small-wavevector mode. Hence, the maximum growth rate (orange squares in Figure 4C) overlaps with the growth rate of the small-wavevector mode (orange triangles) for $D\gtrsim 10^{-2}$. Remarkably, while for a linear source profile, a sufficiently high diffusion coefficient could bring down the maximal perturbation growth rate arbitrarily close to zero (Figure 2C, Eq. (19)), this is not possible any more for a localized source. Instead, for a localized source, the maximum perturbation growth rate is always larger than $\approx 2$ in units of the free deformation rate (Figure 4C).
Figure 5: We introduce co-deforming coordinates to study a deforming system, Eqs. (23) and (24). (A) In lab-frame coordinates, a localized source $s(\bm{r},t)$ is not stationary when the system is deforming affinely, because its width changes proportionally with the system dimension $L_{x}$ due to advection. (B) However, in co-deforming coordinates, the localized source $s(\bm{\mathrm{bar}{r}},t)$ remains stationary under affine deformation.
V Freely deforming system
Here we study the linear stability of a system with a localized source $s_{0}$, Eq. (20), under free deformation. In this case, even for homogeneous shear deformation, the source profile $s_{0}$ is not stationary, but becomes distorted due to advection with the pure shear deformation (Figure 5A).
However, the source profile is stationary with respect to “co-deforming” coordinates. Following our earlier work [32], we define the co-deforming coordinates, $\bm{\mathrm{bar}{r}}=(\mathrm{bar}{x},\mathrm{bar}{y})$, based on the lab-frame coordinates $\bm{r}=(x,y)$ by rescaling both axes:
$\mathrm{bar}{x}$ $:=\frac{L_{x}(0)}{L_{x}(t)}x=l_{x}^{-1}(t)x$ (23)
$\mathrm{bar}{y}$ $:=\frac{L_{y}(0)}{L_{y}(t)}y=l_{x}(t)y,$ (24)
where $l_{x}(t):=L_{x}(t)/L_{x}(0)$ is the box shear. Moreover, we also introduce a flow field in co-deforming coordinates, $\bm{\mathrm{bar}{{\mathrm{v}}}}(\bm{\mathrm{bar}{r}})$, which corresponds to non-affine flows that may occur in addition to the pure shear deformation of the periodic box (Appendix A.3). Thus, for $\bm{\mathrm{bar}{{\mathrm{v}}}}=\bm{0}$ the system deforms affinely. In this case, the source field $s_{0}(\bm{\mathrm{bar}{r}})$ is stationary with respect to the co-deforming coordinates (Figure 5B).
V.1 No morphogen diffusion, $D=0$
Without diffusion, the morphogen concentration given by $c_{0}(\mathrm{bar}{x})=s_{0}(\mathrm{bar}{x})$ is stationary with respect to the co-deforming coordinates, and the system deforms affinely, $\bm{\mathrm{bar}{{\mathrm{v}}}}_{0}=\bm{0}$. This is again due to the homogeneity of $c_{0}$ (and thus the active stresses) in $y$ direction combined with incompressibility (Appendix B.2).
Figure 6: Stability of a freely deforming system with a non-diffusive morphogen, $D=0$. (A) Time evolution of aspect ratios of a gradient-extensile system (green) and a gradient-contractile system (orange) starting from $L_{x}(0)/L_{y}(0)=0.04$ and $L_{x}(0)/L_{y}(0)=25$, respectively. (B) The corresponding absolute values of the shear rates of the systems in panel A. (C) Instantaneous perturbation growth rates over time for the systems in panel A. Square symbols display the maximal growth rates $\omega_{\mathrm{max}}$, whereas triangles display growth rates of small-wavenumber perturbations $\omega_{\mathrm{s}}$ (see Figure 3E). (D) Same as in panel C, where the perturbation growth rates are rescaled by the instantaneous shear rate.
System dimension dynamics. The dynamics of system dimension $L_{x}$, and thus box shear $l_{x}$, is given by Eq. (7). Thus, in the stationary state, the shear rate is essentially given by the integral over $(\partial_{x}c_{0})^{2}$. Because the morphogen profile, $c_{0}$, is stationary in co-deforming coordinates, we have that for varying system size, $(\partial_{x}c_{0})^{2}\sim 1/l_{x}^{2}$. So, using dimensionless units (subsection II.3), we can write
$\frac{1}{l_{x}}\,\frac{\mathrm{d}l_{x}}{\mathrm{d}t}=-\frac{\mathrm{sgn}(\alpha)}{l_{x}^{2}L_{x}^{2}(0)}.$ (25)
The solution of this is:
$l_{x}(t)=\sqrt{1-\frac{2\mathrm{sgn}(\alpha)t}{L_{x}(0)^{2}}}.$ (26)
Thus, gradient-extensile systems ($\alpha<0$) deform slower and slower due to the widening of the morphogen profile (green curves in Figure 6A,B). Meanwhile, gradient-contractile systems ($\alpha>0$) deform faster and faster over time due to the compression of the morphogen profile, until the deformation rate diverges at some critical time point $t_{\mathrm{crit}}=L_{x}^{2}(0)/2$ (orange curves in Figure 6A,B).
Perturbation growth rates. Without diffusion, the dynamics of source and morphogen perturbations, $\delta s$ and $\delta c$, are independent of the box shear perturbation, $\delta l_{x}$ (Eqs. (50) and (51) in Appendix C.1), but they still do depend on box shear $l_{x}(t)$ itself. Thus, to characterize the system’s stability over time, we can use our linear stability analysis for the fixed-size system (section IV and Figure 3), were we use the time-dependent $l_{x}$ according to Eq. (26). The results are shown in Figure 6C: As in section IV and Figure 3, the gradient-contractile systems are always unstable, where even their smallest-wavevector perturbation grows (orange symbols), while the gradient-extensile system is always marginally stable with a decaying smallest-wavevector perturbation (green triangles). We moreover observe that for gradient-contractile systems, the perturbation growth rates become substantially larger at late times, where $l_{x}$ is small. Meanwhile, for gradient-extensile systems, the magnitude of the decay rates is substantially larger at earlier times, where again $l_{x}$ is small. This is likely due to the stronger localization of the source region due to the smaller $l_{x}$; indeed, in Figure 3I we showed that the magnitudes of the perturbation growth rates are larger for smaller source sizes $\mathrm{{w}}$. To better compare to the box shear rate, in Figure 6D we normalize these perturbation growth rates by the instantaneous box shear rate from Figure 6B. We find that the perturbation growth rate magnitudes normalized by the instantaneous shear rates vary much less over time, and is typically on the order of $10$, both for gradient-contractile and -extensile systems. They only become a little smaller for systems with small $l_{x}$ (i.e. late times for gradient-contractile systems, and early times for gradient-extensile systems). Figure 6D allows to predict the expected perturbation growth rates for a biological system of known instantaneous free deformation rate.
V.2 Finite morphogen diffusion, $D>0$
In a fixed-size system with the source field $s_{0}$ given by Eq. (20) and with $D>0$, the interaction between production, diffusion, and degradation allowed for existence of a stationary state (section IV). However, when such a system is allowed to freely deform, it does not reach a stationary state – not even in co-deforming coordinates. This is because diffusion as observed in co-deforming coordinates depends on box shear $l_{x}$ through the transformation rules in Eqs. (23) and (24). Thus, as a consequence of the permanently changing $l_{x}$, there is no stationary state in the freely deforming case.
[…]
Figure 7: Stability of a freely deforming system with a diffusive morphogen, where we set $D=10^{-2}$ and in the limit $\mathrm{{w}}\rightarrow\infty$, where the source has a cosine shape. (A) Aspect ratio over time for the gradient-extensile (green) and contractile (orange) systems, starting at initial aspect ratios of $L_{x}(0)/L_{y}(0)=1/225$ and $L_{x}(0)/L_{y}(0)=25$, respectively. (B) The corresponding absolute values of the shear rates of the systems in panel A. Inset: zoomed around the sharp drop in the gradient-contractile case. (C) Illustration of the affine diffusion-shear instability: diffusion reduces the amplitude $\hat{c}$ of the morphogen field, but a larger $L_{x}$ decreases the effectiveness of diffusion. Thus, a larger $L_{x}$ leads to a larger $\hat{c}$, which in turn increases active shear and thus $L_{x}$. (D) Local Lyapunov exponents (LLE) for the gradient-extensile system over time. Green crosses indicate the numerically computed LLE of the affine diffusion-shear perturbation with the largest growth rate, and the blue curve displays the analytical prediction according to Eq. (88) in Appendix F.2.1. Green squares denote the maximal LLE of all other perturbations. (E) The same results as in panel D, but for the gradient-contractile system.
System size dynamics. In a gradient-extensile system, we observe that the aspect ratio increases over time as expected (green curve in Figure 7A). However, the shear rate first increases and then decreases (green curve in Figure 7B), in contrast to the case without diffusion, where it monotonically decreases (Figure 6B). For gradient-contractile systems, we observe a much more erratic system size behavior (orange curves in Figure 7A,B). In particular, the shear rate shows a transient, but pronounced peak before quickly switching to much slower dynamics (Figure 7B inset). We discuss the reason for these $l_{x}$ dynamics in detail using scaling arguments in Appendix F.1.
[…]
We first briefly discuss the local Lyapunov exponents in the limit of no diffusion, $D=0$ (previous section, subsection V.1). For $D=0$, the linearized dynamics of source and morphogen, Eqs. (50) and (51) (Appendix C.1), do not depend on the perturbation of the box shear, $\delta l_{x}$. Thus, the perturbation dynamics of $s$ and $c$ decouples from that of $l_{x}$, and most of the Lyapunov spectrum is what we already discussed in the previous section. The remaining eigenvalue of the local Lyapunov spectrum is given by the system size perturbation dynamics:
$\frac{\mathrm{d}\delta l_{x}}{\mathrm{d}t}=\frac{\mathrm{sgn}(\alpha)}{l_{x}^{2}(t)L_{x}^{2}(0)}\delta l_{x}.$ (27)
The prefactor in this equation is positive for a gradient-contractile system and negative for a gradient-extensile system. Hence, for $D=0$, the gradient-contractile system has positive local Lyapunov exponents, i.e. the system dynamics strongly depends on the initial conditions. Meanwhile for gradient-extensile systems, the maximal local Lyapunov exponent is zero, and so the system dynamics depends less strongly on the initial conditions.
[…]
For the gradient-contractile case, we found for $D=0$ several modes with a positive growth rate. For $D>0$ this is still true (orange squares in Figure 7E), and – like in the gradient-extensile case – we again have an eigen mode of the Jacobian that is an affine diffusion-shear perturbation, i.e. it does not affect source field and non-affine flows (black crosses in Figure 7E compare predicted growth rate marked by the blue line).
Finally, we studied how the instability due to the affine diffusion-shear perturbation is changed for the general case of finite source width $\mathrm{{w}}$, and we find that this perturbation has a positive growth rate whenever the order of magnitude of the diffusion length becomes larger than the source width, $\lambda\gtrsim\mathrm{{w}}$ (Appendix F.2.2).
VI Discussion
In developing animals, biological tissues can undergo active anisotropic deformation [1]. However, classical active matter results predict that such deformations are unstable under certain conditions [24, 25, 50]. In several systems, the local direction of active tissue deformation is controlled by the gradient of a protein concentration field [12, 30, 7]. Such proteins are called morphogens. They are typically produced by the cells in a localized source region of the tissue, and they can form gradients through diffusion and degrade [1]. However, both morphogen pattern and source region are also advected by the flows generated by the active deformation [37]. Here, we studied the stability of a system which includes source field, morphogen field, and morphogen-gradient-controlled active tissue deformation. Similar to what we found earlier without explicit source region [32], we find that systems where the active stresses are gradient-extensile can be marginally stable, while gradient-contractile systems are always unstable. This is consistent with an observed abundance of gradient-extensile systems in the biological literature [5, 29, 51, 3, 44, 7, 41, 53, 54, 46] and the virtual absence of gradient-contractile systems (also discussed in [55]), suggesting that the instability discussed here may act as an evolutionary selection criterion.
We further find that in the gradient-contractile case, neither degradation nor diffusion can stabilize the system. Degradation only acts – in conjunction with production – to decrease the difference between morphogen and source fields. Meanwhile the system is still subject to the instability. Also, while diffusion can decrease the growth rates of perturbations in the gradient-contractile case, we find that it never suffices to fully stabilize the system. This is because both morphogen and source fields advect with the active flows, while only the morphogen diffuses. In the case of a linear source profile, the maximal perturbation growth rate could become arbitrarily small for sufficiently large diffusion coefficient. However, in the case of a localized source, the maximal perturbation growth rate is always at least $\approx 2$ times the free deformation rate.
We also identify another instability that is not related to the classical active matter instability, which we call “affine diffusion-shear instability”, and which also appears in the gradient-extensile case. More precisely, this instability corresponds to a positive local Lyapunov exponent of a dynamically changing state. It arises from the interplay between the pure shear of the system and the amplitude of the morphogen profile, and in tissues with a localized source whose width is much smaller than the morphogen diffusion length $\sqrt{D/k_{\mathrm{d}}}$. In this case, the extent of the morphogen profile is defined by the diffusion length, while its amplitude is defined by the width of the source region. As a consequence, when the system shear is perturbed to be slightly larger, and consequentially the source width increases, thus the amplitude of the morphogen field increases, which promotes a faster shear. In the context of deforming biological tissues, a positive Lyapunov exponent means that the time-dependent deformation of such tissues is expected to be not very reproducible. In many developing tissues, the morphogen profile has a larger extent than the source region [1, 33, 5], and our work suggests that this instability should occur. This raises the question of how this instability is prevented in developing systems. One possibility might be that there is a saturation in the active stresses as a function of the magnitude of the morphogen gradient. Such a saturation could be mediated for instance through cell polarity [32]. Whether this could be sufficient to suppress this instability is an interesting question for future research.
Our study further motivates experiments, and a key question is whether there are at all any gradient-contractile biological systems. There are many systems where morphogen profiles are known to be necessary for active anisotropic tissue deformation. However, it is most often not yet known whether the gradient of the morphogen controls the directionality or only the magnitude of the active stresses. Thus, experiments are needed that disentangle both. For instance, to modify the direction of the morphogen pattern while keeping the overall magnitude the same, one could combine combine null-mutants with a graded over-expression [30] or the external supply of a gradient, e.g. through morphogen-soaked beads [27, 1]. We believe that combining insights from active matter theory with experiments in this way will help decode the biological mechanisms underlying oriented tissue deformation and their robustness.
Acknowledgements.
We thank the Centre Interdisciplinaire de Nanoscience de Marseille (CINaM) for providing office space. The project leading to this publication has received funding from France 2030, the French Government program managed by the French National Research Agency (ANR-16-CONV-0001), and from the Excellence Initiative of Aix-Marseille University - A*MIDEX. This project was also supported by the grant RobustTissue attributed to M.M. by the French National Research Agency (ANR-22-CE30-0039).
Appendix A Co-deforming coordinates
A.1 Definition
Following Ref. [32], we introduce the co-deforming coordinate system, $(\bm{\mathrm{bar}r},\mathrm{bar}t)=(\mathrm{bar}x,\mathrm{bar}y,\mathrm{bar}t)$, to analytically solve the linearized dynamics of systems under pure shear deformation.
The co-deforming coordinates map to the lab coordinates $(\bm{r},t)=(x,y,t)$ in the following way:
$r_{i}$ $=s_{ij}(\mathrm{bar}{t})\mathrm{bar}{r}_{j}$ (28)
$t$ $=\mathrm{bar}{t},$ (29)
where $\bm{s}(\mathrm{bar}{t})$ is a time-dependent shear tensor, given by
$\bm{s}(\mathrm{bar}{t})=\begin{pmatrix}l_{x}(\mathrm{bar}{t})&0\\ 0&l_{x}^{-1}(\mathrm{bar}{t})\end{pmatrix}.$ (30)
Thus, while at some time $t$, lab coordinates range from $0\leq x<L_{x}(t)$ and $0\leq y<L_{y}(t)$, co-deforming coordinates map these affinely to the box dimensions at time zero, with $0\leq\mathrm{bar}{x}<L_{x}(0)$ and $0\leq\mathrm{bar}{y}<L_{y}(0)$.
Note that for fixed system dimensions, where $L_{x}(t)=L_{x}(0)$, and thus $l_{x}(t)=1$, both, co-deforming and lab coordinates coincide, $\bm{\mathrm{bar}{r}}\equiv\bm{r}$.
A.2 Partial derivatives
As a direct consequence of Eqs. (28) and (29), partial derivatives of some quantity $f$ transform as:
$\mathrm{bar}\partial_{j}f$ $:=\frac{\partial f(\bm{\mathrm{bar}{r}},\mathrm{bar}{t})}{\partial\mathrm{bar}{r}_{j}}=(\partial_{i}f)s_{ij}$ (31)
$\mathrm{bar}\partial_{t}f$ $:=\frac{\partial f(\bm{\mathrm{bar}{r}},\mathrm{bar}{t})}{\partial\mathrm{bar}{t}}=\partial_{t}f+(\partial_{i}f)\dot{s}_{ij}\mathrm{bar}{r}_{j},$ (32)
where $\partial_{i}f:=\partial f(\bm{r},t)/\partial r_{i}$, $\partial_{t}f:=\partial f(\bm{r},t)/\partial t$, and $\dot{s}_{ij}:=\mathrm{d}s_{ij}/\mathrm{d}t=\mathrm{d}s_{ij}/\mathrm{d}\mathrm{bar}{t}$. Thus, the partial time derivative in co-deforming coordinates, $\mathrm{bar}\partial_{{t}}f$, i.e. for fixed $\bm{\mathrm{bar}{r}}$, includes a term related to the box shear rate as compared to the partial time derivative with respect to lab coordinates.
A.3 Velocity and velocity gradient
To obtain the mapping for the velocity field, we consider a tracer particle that is perfectly advected with the flows. The velocity of that tracer particle corresponds to a total time derivative ${\mathrm{v}}_{i}=\mathrm{d}r_{i}/\mathrm{d}t$, for which we obtain by insertion of Eq. (28):
${\mathrm{v}}_{i}=\dot{s}_{ij}\mathrm{bar}{r}_{j}+s_{ij}\mathrm{bar}{{\mathrm{v}}}_{j},$ (33)
where $\mathrm{bar}{\mathrm{v}}_{i}:=\mathrm{d}\mathrm{bar}{r}_{i}/\mathrm{d}t=\mathrm{d}\mathrm{bar}{r}_{i}/\mathrm{d}\mathrm{bar}{t}$ is the co-deforming velocity, with $\mathrm{bar}{r}(\mathrm{bar}{t})$ being the tracer trajectory in co-deforming coordinates.
The first term in Eq. (33) corresponds to a motion due to the affine transformation according to box coordinates. Thus, $\mathrm{bar}{\mathrm{v}}_{i}$ can be interpreted as the non-affine component of the flow field. This can also be seen more explicitly by computing the velocity gradient from Eq. (33):
$\partial_{i}{\mathrm{v}}_{j}=v_{ij}^{0}+s^{-1}_{li}s_{jk}\mathrm{bar}\partial_{l}\mathrm{bar}{{\mathrm{v}}}_{k},$ (34)
where $v^{0}_{ij}:=s^{-1}_{ki}\dot{s}_{jk}$ is the average box deformation rate tensor. For $s_{ij}$ as defined in Eq. (30), the box deformation rate tensor is:
$v^{0}_{ij}=\begin{pmatrix}\dot{l}_{x}/l_{x}&0\\ 0&-\dot{l}_{x}/l_{x}\end{pmatrix}=\begin{pmatrix}\dot{L}_{x}/L_{x}&0\\ 0&-\dot{L}_{x}/L_{x}\end{pmatrix}.$ (35)
A.4 Total derivative
To obtain a transformation formula for the convective derivative, we consider again our tracer and the presence of some spatio-temporal field $f$. The convective derivative corresponds to the total derivative of the value of $f$ that the tracer locally sees. Thus, we expect analogous expressions for the convective derivative in both lab and co-deforming systems, $\dot{f}:=\mathrm{d}f/\mathrm{d}t=\mathrm{d}f/\mathrm{d}\mathrm{bar}t$. Indeed, using Eqs. (31), (32), and (33), we obtain:
$\dot{f}=\partial_{t}f+{\mathrm{v}}_{i}(\partial_{i}f)=\mathrm{bar}\partial_{t}f+\mathrm{bar}{\mathrm{v}}_{i}(\mathrm{bar}\partial_{i}f).$ (36)
A.5 Fourier transform
Moreover, we define the co-deforming Fourier transformation of a quantity $f$ such that
$f(\bm{\mathrm{bar}r})=\sum_{\bm{\mathrm{bar}{q}}}f(\bm{\mathrm{bar}q})\,e^{i\bm{\mathrm{bar}q}\cdot\bm{\mathrm{bar}r}},$ (37)
where the sum is over all co-deforming wave vectors $\bm{\mathrm{bar}{q}}=(\mathrm{bar}{q}_{x},\mathrm{bar}{q}_{y})$ with $\mathrm{bar}q_{i}=2\pi n_{i}/L_{i}(0)$, where $n_{i}\in\mathbb{Z}$ and $i\in\{x,y\}$. A co-deforming Fourier mode with wave vector $\bm{\mathrm{bar}q}$ is distorted over time by the overall system deformation:
$q_{x}(\bm{\mathrm{bar}q},t)$ $=l_{x}^{-1}(t)\mathrm{bar}{q}_{x}$ (38)
$q_{y}(\bm{\mathrm{bar}q},t)$ $=l_{x}(t)\mathrm{bar}{q}_{y}.$ (39)
as in Eq. (37), we have the usual derivation rule, where the Fourier transform of $\mathrm{bar}\partial_{j}f(\bm{\mathrm{bar}r},\mathrm{bar}t)$ is $i\mathrm{bar}{k}_{j}f(\bm{\mathrm{bar}q},\mathrm{bar}t)$. From Eqs. (37) and (28) also follows that a given co-deforming Fourier mode with wave vector $\bm{\mathrm{bar}{q}}$ corresponds to a lab frame Fourier mode with wave vector $\bm{q}$ with components
$q_{i}=\mathrm{bar}{q}_{j}s^{-1}_{ji},$ (40)
because then we have $\bm{\mathrm{bar}{q}}\cdot\bm{\mathrm{bar}{r}}=\bm{q}\cdot\bm{r}$.
Appendix B Dynamics in co-deforming coordinates
Using Appendix A, we can express the dimensionless dynamic equations, Eqs. (10)–(13), in co-deforming coordinates:
$\mathrm{bar}\partial_{t}s+\mathrm{bar}{v}_{i}(\mathrm{bar}\partial_{i}s)$ $=0$ (41)
$\mathrm{bar}\partial_{t}c+\mathrm{bar}{v}_{i}(\mathrm{bar}\partial_{i}c)$ $=\big{(}Ds^{-2}_{ij}\mathrm{bar}\partial_{i}\mathrm{bar}\partial_{j}-k_{\mathrm{d}}\big{)}c+k_{\mathrm{d}}s$ (42)
00 0 $=\frac{1}{4}s_{jl}\partial_{i}^{2}\mathrm{bar}{\mathrm{v}}_{l}-\partial_{j}\Pi^{\prime}+\partial_{i}\tilde{\sigma}^{a}_{ij}$ (43)
$s_{il}\partial_{i}\mathrm{bar}{\mathrm{v}}_{l}$ $=0,$ (44)
where $\tilde{\sigma}^{a}_{ij}$ is given by Eq. (6). Eqs. (43) and (44) are still expressed using lab-frame derivatives, because this will allow for a more convenient solution of the non-affine flows in the next section.
Finally, the system size dynamics for the freely deforming system, Eq. (7), becomes:
$\frac{\mathrm{d}l_{x}}{\mathrm{d}t}=-\mathrm{sgn}(\alpha)l_{x}\iint{\Big{[}l_{x}^{-2}(\mathrm{bar}\partial_{x}c)^{2}-l_{x}^{2}(\mathrm{bar}\partial_{y}c)^{2}\Big{]}\;\mathrm{d}\mathrm{bar}{x}\,\mathrm{d}\mathrm{bar}{y}},$ (45)
where the integral is over the box at time point 0, i.e. the integration variables are the co-deforming coordinates $\mathrm{bar}{x}$ and $\mathrm{bar}{y}$.
B.1 Non-affine flows
For deforming systems, we will generally consider the dynamics of source and morphogen fields in co-deforming coordinates, Eqs. (41) and (42). To obtain the non-affine flows, we solve Eqs. (43) and (44) by taking the lab-frame Fourier transform, which yields after some transformations:
$\mathrm{bar}{\mathrm{v}}_{k}=\frac{4is^{-1}_{ki}q_{l}}{q^{2}}\left(\delta_{ij}-\frac{q_{i}q_{j}}{q^{2}}\right)\tilde{\sigma}^{a}_{lj}$ (46)
with $\tilde{\sigma}^{a}_{ij}$ given by Eq. (6).
Note that Eq. (46) is undefined for $\bm{q}=0$. Indeed, any constant homogeneous velocity can always be added to the flow field. Here and in the following, we will ignore this, because it does not affect the physics in any way.
B.2 Absence of non-affine flows for $\partial_{y}c=0$
If the morphogen concentration field $c$ is homogeneous in $y$ direction, i.e. $\partial_{y}c=\partial_{\mathrm{bar}{y}}c=0$, then there are no non-affine flows, i.e. there are no flows in co-deforming coordinates, $\bm{\mathrm{bar}{\mathrm{v}}}=0$. For fixed system dimensions, this corresponds to a complete absence of flows, $\bm{{\mathrm{v}}}=0$.
Formally, the absence of non-affine flows follows from Eq. (46), which links $\bm{\mathrm{bar}{\mathrm{v}}}$ in Fourier space to the stress tensor in Fourier space. From Eq. (6) we obtain for $\partial_{y}c=0$, that $\tilde{\sigma}^{a}_{xy}=0$ everywhere, and $\tilde{\sigma}^{a}_{xx}$ is independent of $y$. Thus, in Fourier space, $\tilde{\sigma}^{a}_{xx}$ is nonzero only for wave vectors $\bm{q}$ that have a vanishing $y$ component, $q_{y}=0$. Insertion in Eq. (46) yields indeed $\bm{\mathrm{bar}{\mathrm{v}}}=0$.
Appendix C Linearized dynamics
C.1 In co-deforming real space coordinates
[…]
$s(\bm{\mathrm{bar}{r}})$ $=s_{0}(\mathrm{bar}{x})+\delta s(\bm{\mathrm{bar}{r}}),$ (47)
$c(\bm{\mathrm{bar}{r}})$ $=c_{0}(\mathrm{bar}{x})+\delta c(\bm{\mathrm{bar}{r}}),$ (48)
$l_{x}$ $=l_{x0}+\delta l_{x}.$ (49)
[…]
Combining Eqs. (47)–(49) with Eqs. (41), (42), and (45) yields the following linearized dynamics:
$\mathrm{bar}\partial_{t}\delta s$ $=-s_{0}^{\prime}\delta\mathrm{bar}{\mathrm{v}}_{x},$ (50)
$\mathrm{bar}\partial_{t}\delta c$ $=-c_{0}^{\prime}\delta\mathrm{bar}{\mathrm{v}}_{x}+k_{\mathrm{d}}(\delta s-\delta c)$
$\qquad\qquad+Ds^{-2}_{ij}\mathrm{bar}\partial_{i}\mathrm{bar}\partial_{j}\delta c-2D(\mathrm{bar}\partial_{x}^{2}c_{0})\frac{\delta l_{x}}{l_{x0}^{3}},$ (51)
$\frac{\mathrm{d}\delta l_{x}(t)}{\mathrm{d}t}$ $=\frac{2\mathrm{sgn}(\alpha)}{l_{x0}}\iint{(\mathrm{bar}\partial_{x}^{2}c_{0})\delta c\,\mathrm{d}\mathrm{bar}x\mathrm{d}\mathrm{bar}y}-\frac{\mathrm{d}l_{x0}}{\mathrm{d}t}\,\frac{\delta l_{x}}{l_{x0}},$ (52)
where we have introduced the notation $s_{0}^{\prime}:=\mathrm{bar}\partial_{x}s_{0}$ and $c_{0}^{\prime}:=\mathrm{bar}\partial_{x}c_{0}$, and we used a partial integration in the linearized $l_{x}$ dynamics.
C.2 In co-deforming Fourier space coordinates
In co-deforming Fourier space, the linearized dynamics read:
$\mathrm{bar}\partial_{t}\delta s$ $=-s_{0}^{\prime}\ast\delta\mathrm{bar}{\mathrm{v}}_{x}$ (53)
$\mathrm{bar}\partial_{t}\delta c$ $=-c_{0}^{\prime}\ast\delta\mathrm{bar}{\mathrm{v}}_{x}+k_{\mathrm{d}}(\delta s-\delta c)$
$\qquad\qquad-Ds^{-2}_{ij}\mathrm{bar}{q}_{i}\mathrm{bar}{q}_{j}\delta c+2D\mathrm{bar}{q}_{x}^{2}c_{0}\frac{\delta l_{x}}{l_{x0}^{3}},$ (54)
$\frac{\mathrm{d}\delta l_{x}(t)}{\mathrm{d}t}$ $=-\frac{2\mathrm{sgn}(\alpha)}{l_{x0}}\sum_{\bm{\mathrm{bar}{q}}}{\mathrm{bar}{q}_{x}^{2}c_{0}^{\dagger}\delta c}-\frac{\mathrm{d}l_{x0}}{\mathrm{d}t}\,\frac{\delta l_{x}}{l_{x0}}.$ (55)
Here $s_{0}^{\prime}$ and $c_{0}^{\prime}$ represent the co-deforming Fourier transforms of $\mathrm{bar}\partial_{x}s_{0}$ and $\mathrm{bar}\partial_{x}c_{0}$, respectively. Furthermore, to rewrite the $\delta l_{x}$ dynamics, we have used Parseval’s identity, where the dagger, $\cdot^{\dagger}$, denotes the complex conjugate. We have moreover used the following definition of convolution operator, $\ast$, in co-deforming Fourier space:
$f(\bm{\mathrm{bar}{q}})\ast g(\bm{\mathrm{bar}{q}}):=\sum_{\bm{\mathrm{bar}{q}^{\prime}}}f(\bm{\mathrm{bar}{q}^{\prime}})g(\bm{\mathrm{bar}{q}}-\bm{\mathrm{bar}{q}^{\prime}}),$ (56)
where the sum is over all co-deforming wave vectors $\bm{\mathrm{bar}{q}^{\prime}}=(\mathrm{bar}{q}^{\prime}_{x},\mathrm{bar}{q}^{\prime}_{y})$ with $\mathrm{bar}q_{i}^{\prime}=2\pi n^{\prime}_{i}/L_{i}(0)$, where $n^{\prime}_{i}\in\mathbb{Z}$ and $i\in\{x,y\}$. For a fixed system size, we ignore the $\delta l_{x}$ dynamics and the $\delta l_{x}$ term in the $\delta c$ dynamics.
To close Eqs. (53)–(55), we still need to insert the $x$-component of non-affine flow field perturbation, $\delta\mathrm{bar}{{\mathrm{v}}}_{x}$. For convenience, we will use lab-frame Fourier modes to compute $\delta\mathrm{bar}{{\mathrm{v}}}_{x}$. Specifically, we use Eq. (46), which expresses $\bm{\mathrm{bar}{\mathrm{v}}}$ in terms of the active stress tensor $\tilde{\sigma}^{a}_{ij}$. We obtain to linear order:
$\delta\mathrm{bar}{\mathrm{v}}_{x}(q,\phi)=\frac{4i\sin{\phi}}{ql_{x0}}\Big{[}\sin{2\phi}\;\delta\tilde{\sigma}^{a}_{xx}-\cos{2\phi}\;\delta\tilde{\sigma}^{a}_{xy}\Big{]},$ (57)
where $q$ and $\phi$ denote the amplitude and orientation, respectively, of the lab-frame wave vector $\bm{q}=q(\cos{\phi},\sin{\phi})$.
We further express $\delta\tilde{\sigma}^{a}_{ij}$ in terms of $\delta c$ using Eq. (6). To linear order, in lab-frame Fourier space:
[…]
The other two components of the symmetric traceless tensor $\delta\tilde{\sigma}^{a}_{ij}$ are obtained through the relations $\delta\tilde{\sigma}^{a}_{yy}=-\delta\tilde{\sigma}^{a}_{xx}$ and $\delta\tilde{\sigma}^{a}_{yx}=\delta\tilde{\sigma}^{a}_{xy}$. The convolution operator in lab-frame Fourier coordinates in Eq. (58) is defined analogously to Eq. (56) as:
$f(\bm{q})\ast g(\bm{q}):=\sum_{\bm{q^{\prime}}}f(\bm{q^{\prime}})g(\bm{q}-\bm{q^{\prime}}),$ (59)
where the sum is over all lab-frame wave vectors $\bm{q^{\prime}}=(q^{\prime}_{x},q^{\prime}_{y})$ with $q_{i}^{\prime}=2\pi n^{\prime}_{i}/L_{i}(t)$, where $n^{\prime}_{i}\in\mathbb{Z}$ and $i\in\{x,y\}$.
Combining (57) and (58), we obtain:
$\begin{split}\delta\mathrm{bar}{\mathrm{v}}_{x}=&-\frac{4\mathrm{sgn}(\alpha)\sin{\phi}}{ql_{x0}^{2}}\bigg{\{}\sin{2\phi}\Big{[}c_{0}^{\prime}\ast(q_{x}\delta c)\Big{]}\\ &\qquad\qquad\qquad\qquad-\cos{2\phi}\Big{[}c_{0}^{\prime}\ast(q_{y}\delta c)\Big{]}\bigg{\}}.\end{split}$ (60)
The system of equations, Eqs. (53)–(55), and (60), is now closed.
C.3 Numerical solution
In the special case of a linear source profile, we can solve the linearized dynamics analytically, because the gradients of $s_{0}$ and $c_{0}$ are constant, and thus the convolutions in Eqs. (53), (54), and (60) collapse to a simple scaling factor. However, in general this is not the case, and we need to solve the linearized dynamics numerically.
To numerically compute the convolutions in Eqs. (53), (54), and (60), we need to take into account that the numerical real-space representations of our fields are spatially discretized. To discuss the main ideas, we will discuss here the case of a single spatial dimension, $\mathrm{bar}{x}$, in co-deforming coordinates. The generalization to lab frame coordinates and/or several dimensions will be straight forward. Specifically, we consider the periodic interval $\mathrm{bar}{x}\in[-L_{x}(0)/2,+L_{x}(0)/2)$ to be discretized by $N_{x}$ equal-length steps $\Delta\mathrm{bar}x=L_{x}(0)/N_{x}$, where $N_{x}$ is an even integer. Any function $f(\mathrm{bar}{x})$ is represented by a set of numbers $f_{k}$ with $k=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$, such that $f_{k}=f(k\Delta\mathrm{bar}x)$. As a consequence, the Fourier series are not infinite any more as in Eq. (37), but instead cut off at some finite wave vectors. In particular, the numerical co-deforming Fourier transform of $f$ is represented by a set of complex numbers $\tilde{f}_{j}$ with $j=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$, where $\tilde{f}_{j}=f(\mathrm{bar}{q}_{x}=j\mathrm{bar}{q}_{0})$ with $\mathrm{bar}{q}_{0}:=2\pi/L_{x}(0)$. Eq. (37) then becomes:
$f_{k}=\sum_{j=-N_{x}/2}^{N_{x}/2-1}{\tilde{f}_{j}e^{2\pi i(jk/N_{x})}},$ (61)
where we used that $\mathrm{bar}{q}_{0}\Delta\mathrm{bar}x=2\pi/N_{x}$. To maintain in this discretized space the property that the Fourier transform of the product of two functions $f$ and $g$ is the convolution of their respective Fourier transforms,
$\widetilde{(fg)}_{j}=(\tilde{f}\ast\tilde{g})_{j},$ (62)
we need to define the following discrete convolution operator:
$(\tilde{f}\ast\tilde{g})_{j}=\sum_{j^{\prime}=-N_{x}/2}^{N_{x}/2-1}{\tilde{f}_{j^{\prime}}\tilde{g}_{j-j^{\prime}}}$ (63)
for $j=-N_{x}/2,-N_{x}/2+1,\dots,0,\dots,N_{x}/2-1$. Yet, we see that the index $j-j^{\prime}$ of $\tilde{g}$ can get outside of the range $-N_{x}/2,-N_{x}/2+1,\dots,N_{x}/2-1$. For Eq. (62) to still be fulfilled, we need to define $\tilde{g}$ to be periodic, such that $\tilde{g}_{(j+nN_{x})}:=\tilde{g}_{j}$ for all $n\in\mathbb{Z}$. In this sense, the convolution in Fourier space of our discretized function is circular. We use such a circular definition to numerically compute the convolutions in Eqs. (53), (54), and (60), to preserve the relation Eq. (62).
To numerically solve the linearized dynamics, Eqs. (53)–(55), we use an exponential Ansatz for the time dependence of the perturbations,
$\delta s(\bm{\mathrm{bar}{q}},t)$ $=\delta\hat{s}(\bm{\mathrm{bar}{q}})e^{\omega t}$ (64)
$\delta c(\bm{\mathrm{bar}{q}},t)$ $=\delta\hat{c}(\bm{\mathrm{bar}{q}})e^{\omega t}$ (65)
$\delta l_{x}(t)$ $=\delta\hat{l}_{x}e^{\omega t}.$ (66)
Inserting this in Eqs. (53)–(55) yields an eigenvalue problem:
[…]
Appendix D Fixed-size system with linear source
For a linear source and morphogen fields, the active flows, Eq. (60), simplify to:
[…]
where $\omega_{\mathrm{act}}$ is the perturbation growth rate for the case of a boundary-provided gradient without source field [32]. With this, the linearized dynamics, Eqs. (53)–(54) become:
$\partial_{t}\delta s$ $=\omega_{\mathrm{act}}\delta c,$ (69)
$\partial_{t}\delta c$ $=\left[\omega_{\mathrm{act}}-Dq^{2}-k_{\mathrm{d}}\right]\delta c+k_{\mathrm{d}}\delta s.$ (70)
Thus, the convolutions in Eqs. (53), (54), and (60) collapse to a constant prefactor. As a consequence, the solutions of the system are Fourier modes. In the following, we first discuss the limiting case of no diffusion before presenting the general solution.
D.1 No diffusion, $D=0$
To solve the eigenvalue problem in this case, we define
$\delta\Delta$ $=\delta c-\delta s$ (71)
$\delta b$ $=\delta c+\frac{k_{\mathrm{d}}}{\omega_{\mathrm{act}}}\delta s.$ (72)
Inserting this into Eqs. (69) and (70), we obtain:
$\partial_{t}\delta\Delta$ $=-k_{\mathrm{d}}\delta\Delta$ (73)
$\partial_{t}\delta b$ $=\omega_{\mathrm{act}}\partial_{t}\delta b.$ (74)
Hence, the difference between both morphogen and source, $\delta\Delta$, decays at rate $k_{\mathrm{d}}$, while a superposition of both, $\delta b$, grows with the same growth rate as the system without source [32].
D.2 General case
With an exponential Ansatz, where $\delta s\sim\delta c\sim e^{\omega t}$, we obtain an eigenvalue problem, whose solutions in $\omega$ are:
$\omega_{\pm}=\frac{1}{2}\left[\mathcal{T}\pm\sqrt{\mathcal{T}^{2}+4k_{\mathrm{d}}\;\omega_{\mathrm{act}}}\right],$ (75)
where $\mathcal{T}=\omega_{\mathrm{act}}-Dq^{2}-k_{\mathrm{d}}$. The corresponding eigenvectors in $(\delta s,\delta c)$ are
$\bm{v}_{+}=\begin{pmatrix}\omega_{+}\\ \omega_{\mathrm{act}}\end{pmatrix}\quad\mathrm{and}\quad\bm{v}_{-}=\begin{pmatrix}\omega_{-}\\ \omega_{\mathrm{act}}\end{pmatrix},$ (76)
respectively.
Appendix E Fixed-size system with localized source
E.1 Value of $s_{b}$ in dimensionless units
As we discussed in subsection II.3, we set the value of $s_{b}$ such that the magnitude of the deformation rate of a system with $L_{x}=L_{y}=1$ is one. This means that we set
$s_{b}=\left[\sum_{q_{x}}{\frac{k_{\mathrm{d}}^{2}|m(q_{x})|^{2}}{(k_{\mathrm{d}}+Dq_{x}^{2})^{2}}}\right]^{-1/2}$ (77)
where $m(q_{x})$ is the Fourier transform of the function
$m(x)=\exp{\bigg{(}\frac{\cos{(q_{0}x)}}{(q_{0}\mathrm{{w}})^{2}}\bigg{)}}$ (78)
which is defined on the interval $x\in[-1/2,1/2)$, and where $q_{0}=2\pi/L_{x}$. Indeed, using Eqs. (7) and (21), one obtains that with the value of $s_{b}$ from Eq. (77), a von Mises source profile, Eq. (20), and $L_{x}=L_{y}=1$, the absolute value of the shear rate is one.
Figure 8: Linear stability of gradient-extensile systems with $\mathrm{{w}}\rightarrow\infty$, i.e., where the source profile converges to a single cosine mode. Green dots and green circles represent the numerical analysis with and without discrete periodic convolution, respectively. Inset showing that perturbation modes with $\omega>0$ disappear when convolution is not periodic.
E.2 Numerically growing modes in gradient-extensile systems
In section IV, we found that in a gradient-extensile systems, some modes showed a positive growth rate. Yet, when we increased the spatial discretization $N_{x}$, the value of the growth rate decreased as a power law with $N_{x}$ (Figure 3H), suggesting that these positive growth rates may just be a numerical artifact. Here we discuss this question further.
We discuss the positive growth rates for the limit $\mathrm{{w}}\rightarrow\infty$, for which the source profile $s_{0}(\bm{r})$, defined in Eq. (20), becomes a cosine. We obtained the modes with positive growth rates, using the approach described in Appendix C.3. Yet, for this we used the circular convolutions, Eq. (63), which in a way is a natural choice, imposed by the fact that we do not have infinitely fine resolution of $s(x)$ and $c(x)$, which would allow to have the information about these fields at arbitrarily large wave number $\bm{q}$. Instead, because these fields are discretized, the Fourier amplitude at high wave vectors just become periodic, as described in Appendix C.3. In other words, due to a limited resolution, we do not know what the amplitude for high wave vectors is and we “just” set them to some values assuming periodicity. One could also use a different approach, by assuming that the Fourier amplitudes at these high wave vectors just vanishes. This corresponds to an alternative numerical implementation of the Fourier-space convolution. Specifically, we replace the circular convolutions by “non-circular” convolutions where we ignore the periodicity of the discrete Fourier transforms, i.e., we replace Eq. (63) by:
$(\tilde{f}\ast\tilde{g})_{j}=\sum_{j^{\prime}=-N_{x}/2}^{N_{x}/2-1}{\tilde{f}_{j^{\prime}}\tilde{g}_{j-j^{\prime}}\Theta_{j-j^{\prime}}}$ (79)
where
$\Theta_{k}=\begin{cases}1&\text{for $-N_{x}/2\leq k\leq N_{x}/2-1$,}\\ 0&\text{otherwise.}\end{cases}$ (80)
The factor $\Theta_{j-j^{\prime}}$ explicitly ensures that there is any periodicity in $\tilde{g}_{j-j^{\prime}}$ is ignored. We compare the solutions of the eigenvalue problem, Eq. (67), between circular and non-circular convolution in Figure 8, and we find that while most of the modes did not change much, the modes with positive growth rates disappeared when using the non-circular convolutions. This further suggests that these positive growth rates are merely an artifact of our spatial discretization.
Figure 9: System dynamics of a freely deforming system with a diffusive morphogen, $\tilde{v}_{xx}^{0}=\tilde{v}_{xx}^{\mathrm{free}},D=10^{-2}$. (A) Morphogen magnitude over time for the gradient-extensile case, where we compare the dynamic system (green) with the stationary value (Eq. (83), dashed red) as a function of time (B) Absolute values of shear rates of the systems in panel A. (C,D) Morphogen and shear rate magnitudes for the gradient-contractile case. (E) Decomposition of the partial time derivative of the dynamic morphogen amplitude in the gradient-contractile case (see text).
Appendix F Deforming system with localized source
F.1 Discussion of the system size dynamics
Starting with a gradient-extensile system with dimension $L_{x}=1/15$, we observe that the aspect ratio increases over time as expected (green curve in Figure 7A). However, the shear rate first increases and then decreases (green curve in Figure 7B), in contrast to the case without diffusion, where it monotonically decreases (Figure 6B). To understand this behavior, we first note that without any non-affine flows (i.e. $\bm{\mathrm{bar}{{\mathrm{v}}}}=\bm{0}$), both morphogen and source profiles keep their cosine profiles, $s(\bm{r},t)=\hat{s}[\cos{(2\pi x/L_{x})}+1]$ and $c(\bm{r},t)=\hat{c}(t)\cos{(2\pi x/L_{x})}+c_{\mathrm{offset}}(t)$, while only the morphogen amplitude $\hat{c}(t)$ and offset $c_{\mathrm{offset}}(t)$ vary over time. Without non-affine flows we can thus simplify the system dynamics to:
$\frac{\mathrm{d}\hat{c}}{\mathrm{d}t}$ $=-k_{\mathrm{d}}(\hat{c}-\hat{s})-4\pi^{2}D\frac{\hat{c}}{L_{x}^{2}}$ (81)
$\frac{\mathrm{d}L_{x}}{\mathrm{d}t}$ $=-\frac{2\pi^{2}\mathrm{sgn}(\alpha)\hat{c}^{2}}{L_{x}},$ (82)
where the offset $c_{\mathrm{offset}}(t)$ does not matter for the system size dynamics. We also noted that at each time point the morphogen amplitude $\hat{c}$ is relatively well approximated by its stationary value $\hat{c}_{\mathrm{s}}$ for given $L_{x}$ (Figure 9A). This corresponds to the adiabatic limit where the $c$ dynamics is fast as compared to the $L_{x}$ dynamics. This stationary value $\hat{c}_{\mathrm{s}}$ is given by Eq. (81) for $\mathrm{d}\hat{c}/\mathrm{d}t=0$, which implies:
$\hat{c}_{\mathrm{s}}=\left[1+\frac{4\pi^{2}D}{k_{\mathrm{d}}L_{x}^{2}}\right]^{-1}\hat{s}.$ (83)
Indeed, when numerically solving the system size dynamics given by Eq. (82) with the stationary solution for $\hat{c}_{\mathrm{s}}$, Eq. (83) instead of the actual $\hat{c}$ dynamics, Eq. (81), we get a similar behavior (Figure 9B). This behavior can be readily explained as initially $L_{x}$ is small, and thus the diffusion term in $\hat{c}$ dominates, such that $\hat{c}_{s}\sim L_{x}^{2}$. Inserting this into Eq. (82) gives a shear rate scaling of $(\mathrm{d}L_{x}/\mathrm{d}t)/L_{x}\sim L_{x}^{2}$, which leads for a gradient-extensile system ($\mathrm{sgn}(\alpha)=-1$) to an accelerating pure shear. Intuitively, this is because for small $L_{x}$, diffusion dominates, which generally lowers the amplitude of the morphogen field. In this limit, increasing $L_{x}$ decreases diffusion and thus increases the morphogen amplitude and, as a consequence, the shear rate. However, as soon as $L_{x}$ becomes large enough, the diffusion term in Eq. (83) can be neglected, and we have almost constant $\hat{c}$, such that the shear rate scales as $(\mathrm{d}L_{x}/\mathrm{d}t)/L_{x}\sim L_{x}^{-2}$. Hence, the shear decelerates over time due to the expansion of the morphogen field (subsection V.1).
For gradient-contractile systems, we observe a much more erratic system size behavior. The orange curves in Figure 7A,B show example trajectories for a system with initially $L_{x}=5$. In particular, the shear rate shows a transient, but pronounced peak before quickly switching to much slower dynamics (Figure 7B inset). This behavior can be quantitatively understood in terms of three distinct phases. In a first phase, $L_{x}$ is large, such that the diffusion term can be neglected in the morphogen dynamics, so that, according to Eq. (81), $\hat{c}$ keeps its initial value of approximately $\hat{s}$. Thus, according to Eq. (82), the shear rate initially scales as $(\mathrm{d}L_{x}/\mathrm{d}t)/L_{x}\sim-L_{x}^{-2}$, leading to a decrease in $L_{x}$ and an increase in the shear rate, as in the case without diffusion (subsection V.1). The second phase starts when $L_{x}$ becomes small enough for the diffusion term to dominate in the $\hat{c}$ dynamics, Eq. (81), such that $\mathrm{d}\hat{c}/\mathrm{d}t=-4\pi^{2}D\hat{c}/L_{x}^{2}$ (Figure 9E). Dividing Eq. (82) by $\mathrm{d}\hat{c}/\mathrm{d}t$, we get $\mathrm{d}L_{x}/\mathrm{d}\hat{c}=\mathrm{sgn}(\alpha)\hat{c}L_{x}/2D$, which implies that $L_{x}\sim\exp{[\mathrm{sgn}(\alpha)\hat{c}^{2}/4D]}$. This limit corresponds to the transient peak in the shear rate (Figure 9C,D). During this transient phase, $\hat{c}$ is rapidly decreasing, while for very small $\hat{c}$ the system dimension $L_{x}\sim\exp{[\mathrm{sgn}(\alpha)\hat{c}^{2}/4D]}$ varies only little as compared to $\hat{c}$. Thus, as $\hat{c}$ decreases, the diffusion term in Eq. (81) at some point becomes smaller than the source term $k_{\mathrm{d}}\hat{s}$, at which point this transient stops. During the subsequent third phase, the shear rate becomes small due to the small $\hat{c}$, which allows the morphogen dynamics to be close to its stationary state. But since $L_{x}$ is also small, the diffusion term still dominates in Eq. (83), and so the morphogen amplitude scales as $\hat{c}\sim L_{x}^{2}$, such that the shear rate scales as $(\mathrm{d}L_{x}/\mathrm{d}t)/L_{x}\sim L_{x}^{2}$, and thus slows down over time. Taken together, while the shear rate varies for both, gradient-extensile and gradient-contractile systems, its behavior is much more erratic in the gradient-contractile case.
F.2 Affine diffusion-shear instability
To discuss the affine diffusion-shear instability, we start from the linearized dynamics in Fourier space, Eqs. (53)–(55), where we set non-affine flows to zero, $\bm{\delta\mathrm{bar}{\mathrm{v}}}=0$, we ignore any variations of the source field, $\delta s=0$, and we ignore any $y$-dependence of the perturbations. We obtain from Eqs. (54) and (55) with an exponential Ansatz for the time dependency, $\delta c(\mathrm{bar}q_{x},t)=\delta\tilde{c}(\mathrm{bar}q_{x})e^{\omega t}$ and $\delta l_{x}(t)=\delta\tilde{l}_{x}e^{\omega t}$:
$\omega\delta\tilde{c}(\mathrm{bar}q_{x})$ $=-k_{d}\big{(}1+\lambda^{2}q_{x}^{2}\big{)}\delta\tilde{c}(\mathrm{bar}q_{x})+2k_{d}\lambda^{2}q_{x}^{2}c_{0}(\mathrm{bar}q_{x})\frac{\delta\tilde{l}_{x}}{l_{x}}$ (84)
$\omega\frac{\delta\tilde{l}_{x}}{l_{x}}$ $=-2\mathrm{sgn}(\alpha)\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}c_{0}^{\dagger}(\mathrm{bar}q_{x})\delta\tilde{c}(\mathrm{bar}q_{x})}$
$\qquad\qquad+\mathrm{sgn}(\alpha)\frac{\delta\tilde{l}_{x}}{l_{x}}\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}.$ (85)
Here, we also substituted the diffusion length scale $\lambda=\sqrt{D/k_{d}}$ and the lab-frame wave vector $q_{x}=\mathrm{bar}q_{x}/l_{x}$. Iff this linear system of equations has a positive eigenvalue $\omega$, then there is an instability.
F.2.1 Cosine source profile
For the case of a cosine profile, the only nonzero values of $c_{0}(\mathrm{bar}q_{x})$ are $c_{0}\equiv c_{0}(\pm\mathrm{bar}q_{0})$ with $\mathrm{bar}q_{0}=2\pi/L_{x}(0)$. Then, Eqs. (84) and (85) become:
$\omega\delta\tilde{c}$ $=-k_{d}\big{(}1+\lambda^{2}q_{0}^{2}\big{)}\delta\tilde{c}+2k_{d}\lambda^{2}q_{0}^{2}c_{0}\frac{\delta\tilde{l}_{x}}{l_{x}}$ (86)
$\omega\frac{\delta\tilde{l}_{x}}{l_{x}}$ $=-4\mathrm{sgn}(\alpha)q_{0}^{2}c_{0}^{\dagger}\delta\tilde{c}+2\mathrm{sgn}(\alpha)q_{0}^{2}|c_{0}|^{2}\frac{\delta\tilde{l}_{x}}{l_{x}},$ (87)
where $q_{0}(t):=\mathrm{bar}q_{0}/l_{x}(t)=2\pi/L_{x}(t)$. There are two eigenvalues, $\omega_{1}$ and $\omega_{2}$ with $\omega_{1}\leq\omega_{2}$, and the larger eigenvalue is given by: The larger of the two eigenvalues is:
$\omega_{2}=\frac{1}{2}\left[\mathcal{T}+\sqrt{\mathcal{T}^{2}-4\mathcal{D}}\right],$ (88)
where $\mathcal{T}:=-k_{d}(1+\lambda^{2}q_{0}^{2})+2\mathrm{sgn}(\alpha)q_{0}^{2}|c_{0}|^{2}$ and $\mathcal{D}:=-2\mathrm{sgn}(\alpha)k_{d}q_{0}^{2}|c_{0}|^{2}(1-3\lambda^{2}q_{0}^{2})$ are trace and determinant, respectively, of the matrix on the right-hand side of Eqs. (86) and (87).
In the gradient-extensile case, $\mathrm{sgn}(\alpha)=-1$, we have that $\omega_{1}+\omega_{2}=\mathcal{T}<0$. Thus, at least one eigenvalue is negative. The product of both eigenvalues is given by $\omega_{1}\omega_{2}=\mathcal{D}=2k_{d}q_{0}^{2}|c_{0}|^{2}(1-3\lambda^{2}q_{0}^{2})$. Hence, there is a positive eigenvalue iff $13\lambda^{2}q_{0}^{2}>$ 1, i.e. iff $2\pi\lambda\sqrt{3}>L_{x}$.
F.2.2 Arbitrary source profile
For the case of an arbitrary source profile $s_{0}$, we focus on the gradient-extensile case, $\mathrm{sgn}(\alpha)=-1$. We discuss the existence of positive solutions $\omega$ by first expressing $\delta\tilde{c}(\mathrm{bar}q_{x})$ in terms of $\delta\tilde{l}_{x}$ and $\omega$ using Eq. (84):
$\delta\tilde{c}(\mathrm{bar}q_{x})=\frac{2\lambda^{2}q_{x}^{2}c_{0}(\mathrm{bar}q_{x})}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}\,\frac{\delta\tilde{l}_{x}}{l_{x}}.$ (89)
Insertion into Eq. (85) yields:
$\begin{split}\omega&=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{4\lambda^{2}q_{x}^{2}}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}}-\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}\\ &=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{d}}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}}.\end{split}$ (90)
To solving this self-consistent equation for $\omega$, we look for the zeros of the function
$f_{\lambda}(\omega):=-\omega+\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{d}}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}}.$ (91)
Specifically, we now show that iff $f_{\lambda}(0)>0$, there exits a positive solution in $\omega$, i.e. an $\omega>0$ for which $f_{\lambda}(\omega)=0$. This means that there is an affine diffusion-shear instability iff $f_{\lambda}(0)>0$. To show this, we first assume that there is an $\omega>0$ for which $f_{\lambda}(\omega)=0$. Then, we have:
$\begin{split}0&<\omega\\ &\quad=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{d}}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}}\\ &\quad<\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1}{1+\lambda^{2}q_{x}^{2}}}\\ &\qquad=f_{\lambda}(0).\end{split}$ (92)
In these transformations, we have subsequently used $\omega>0$, $f_{\lambda}(\omega)=0$, $\omega>0$, and the definition of $f_{\lambda}(\omega)$. Thus, if there is an $\omega>0$ for which $f_{\lambda}(\omega)=0$, then $f_{\lambda}(0)>0$.
Second, we assume that $f_{\lambda}(0)>0$. To show that there is a positive zero of $f_{\lambda}$, we need to show that there is a positive $\hat{\omega}>0$ for which $f_{\lambda}(\hat{\omega})<0$. Then, the Zwischenwertsatz implies that there is an $\omega$ with $0<\omega<\hat{\omega}$ such that $f_{\lambda}(\omega)=0$. Indeed, for $\hat{\omega}=3\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}$, we find that:
$\begin{split}f_{\lambda}(\hat{\omega})&=-3\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}\\ &\qquad\qquad\qquad+\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\hat{\omega}/k_{d}}{1+\hat{\omega}/k_{d}+\lambda^{2}q_{x}^{2}}}\\ &=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{-2(1+\hat{\omega}/k_{d})}{1+\hat{\omega}/k_{d}+\lambda^{2}q_{x}^{2}}}\\ &<0.\end{split}$ (93)
Thus, if $f_{\lambda}(0)>0$ then $f_{\lambda}$ has a positive zero and the affine diffusion-shear instability exists. Taken together, the sign of $f_{\lambda}(0)$ directly indicates the existence of an instability.
Finally, we discuss the sign of $f_{\lambda}(0)$. Using Eq. (21), we can express $f_{\lambda}(0)$ in terms of the source profile:
$f_{\lambda}(0)=\frac{1}{\lambda^{2}}\sum_{\mathrm{bar}q_{x}}{|s_{0}(\mathrm{bar}q_{x})|^{2}\frac{\lambda^{2}q_{x}^{2}(3\lambda^{2}q_{x}^{2}-1)}{(1+\lambda^{2}q_{x}^{2})^{3}}}.$ (94)
Choosing for $s_{0}$ the von Mises profile in Eq. (20) with finite width $\mathrm{{w}}$, we numerically obtain $f_{\lambda}(0)>0$ iff $\lambda>g\mathrm{{w}}$ with $g\approx 0.639$ in the limit where $\mathrm{{w}}\ll L_{x}$ (i.e. where the von Mises profile becomes a Gaussian with standard deviation $\mathrm{{w}}$). However, the precise value of $g$ depends on the shape of the profile $s_{0}$. For a rectangular profile of $s_{0}$, we numerically even found that the instability always occurs.